Solve the system using the substitution method
0.1x+0.3y=1.6
0.8x-0.7y=5.8

Answer by lwsshak3(11628)

Solve the system using the substitution method
0.1x+0.3y=1.6
0.8x-0.7y=5.8
**
multiply both equations by 10 to remove decimals (move decimal point 1 place right)
x+3y=16
8x-7y=58
..
x+3y=16
x=16-3y
sub (16-3y) for x in second equation
8x-7y=58
8(16-3y)-7y=58
128-24y-7y=58
31y=70
y=70/31
..
x=16-3y
=16-3(70/31)
=16-210/31
=(31*16-210)/31
x=286/31